OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 5n + 6 = -4

Parth (parthkohli):

\(\Large \color{purple}{ 5n = -4 - 6 }\) \(\Large \color{purple}{ 5n = -10 }\) \(\Large \color{purple}{ n = {-10 \over 5} }\) \(\Large \color{purple}{ n = -2 }\)

Parth (parthkohli):

\(\Large \color{purple}{5(-2) + 6 = -4 }\) \(\Large \color{purple}{ -10 + 6 = -4 }\) \(\Large \color{purple}{ -4 = -4 }\) Checked :D

Parth (parthkohli):

\(\Large \color{blue}{\text{Got it?} }\)

OpenStudy (kropot72):

5n + 6 = -4 Subtract 6 from both sides of the equation: 5n + 6 -6 = -4 -6 5n = -10 Divide both sides by 5: 5n/5 = -10/5 n = -2

OpenStudy (anonymous):

yeah totally thanks :)

OpenStudy (kropot72):

You're welcome :)